OpenStudy (anonymous):

Because I want to solve an equation for myself, I would just like to know how to change an equation into slope-intercept form?

OpenStudy (anonymous):

depends on what form it is currently

OpenStudy (anonymous):

if it looks like \[ax+by=c\] then solve for \(y\) via \[by=-ax+c\\ y=-\frac{a}{b}+\frac{c}{b}\]

OpenStudy (anonymous):

what the hell was that?!!

OpenStudy (anonymous):

if the equation is in point slope form like \[y-y_1=m(x-x_1)\] then distribute first \[y-y_1=mx-mx_1\] then add \(y_1\)

OpenStudy (anonymous):

my equation looks like ax+by=c